OpenStudy (dan815):

Fun question... Three people are sitting at a restaurant table, a grandfather, his son and his grandson.The waiter asked the boss what are their ages, the boss replied "The product of their ages is 4032, the sum of their ages is twice my age and if you divide the grandfather's age by the grandson's age you get twice your age". What is the age of the son? All ages are integer values

OpenStudy (smored):

what is the boss's age and the waiter's age?

OpenStudy (smored):

"the sum of their ages is twice my age (boss) and if you divide the grandfather's age by the grandson's age you get twice your age (waiter)"

OpenStudy (zzr0ck3r):

3 equations, 5 unknowns....

ganeshie8 (ganeshie8):

k : grand son(kid) s : son g : grand father b : boss w : waiter ksg = 4032 k+s+g = 2b g/k = w

ganeshie8 (ganeshie8):

solve in naturals

OpenStudy (zzr0ck3r):

g/k = 2w

ganeshie8 (ganeshie8):

i would start here : http://www.wolframalpha.com/input/?i=prime+factorization+of+4032

OpenStudy (smored):

k, s, and g could be any three #'s that are factors of 4032 because b and w are not specified.

ganeshie8 (ganeshie8):

ohk good catch :) corrected below : ksg = 4032 k+s+g = 2b g/k = 2w

OpenStudy (smored):

let me correct myself, k, s, and g can be any 3 #'s that multiply to be 4032

ganeshie8 (ganeshie8):

also one more constrain : k < s < g

ganeshie8 (ganeshie8):

so, the system is : ksg = 4032 k+s+g = 2b g/k = 2w k < s < g all naturals

OpenStudy (ranga):

nothing funny like the grandfather is the boss or some such thing?

OpenStudy (smored):

there's more than one answer, since the b and w are not specified.

ganeshie8 (ganeshie8):

\(g/k = 2w \implies g = 2kw\) that gives \(g \) must be even

OpenStudy (dan815):

solve it!! ill give u a hint dont pay too much attention to what age is it really possible to have kids and that stuff lol

OpenStudy (dan815):

the numbers were chosen to just to make it funny i think

OpenStudy (dan815):

just to mess with all those guessers haha

OpenStudy (dan815):

you are on right path!! ganeshie8

ganeshie8 (ganeshie8):

the kid cannot have 7 as a factor : \(k = 2^m3^n\) \(0\le m \le 2, ~~0\le n \le 1\)

ganeshie8 (ganeshie8):

m and n cannot be 0 at the same time, so that gives 5 cases to consider

ganeshie8 (ganeshie8):

oh they can be 0 at the same time as well, then the kid will be 1 year old

ganeshie8 (ganeshie8):

\(k = 1\) seem to fit all constraints ?

OpenStudy (dan815):

the question is actually pretty deep, its a number theory question, but the math isnt complicated its just tricky

OpenStudy (dan815):

u gotta start here first

OpenStudy (dan815):

xyz=4032 x/z=2W x+y+z=2B x=2wz 2wZ^2*Y=4032 Z=sqrt(2016/yw)

OpenStudy (dan815):

try to take it from there

OpenStudy (dan815):

gota do some number theory deductions from there

ganeshie8 (ganeshie8):

beautiful ! yw can be : 2*7 or 2^3*7 or 2^5*7 or 2*3^2*7 or 2^3*3^2*7 or 2^5*3^2*7

ganeshie8 (ganeshie8):

kid = {12, 6, 3, 4, 2, 1}

ganeshie8 (ganeshie8):

it can be any of above 6... thats it right /

OpenStudy (dan815):

theerse more

OpenStudy (dan815):

theres actually a unique answer

OpenStudy (dan815):

i didnt give u much it was just to get started, i dont wanna ruin it for u lol

ganeshie8 (ganeshie8):

oh we dint fold in below constraint yet x+y+z=2B

ganeshie8 (ganeshie8):

that constraint actually satisfies all the 6 values for kid acutally

ganeshie8 (ganeshie8):

\(\color{red}{\#\# k=1, s = 3^2, g = 2^67}\) \(ksg = 2^63^27 = 4032\) \(k+s+g = 2(5 + 2^57)\) = 2b \(g/k = 2(2^47) = 2w\) \(\color{red}{\#\# k=2, s = 23^2, g = 2^47}\) \(ksg = 2^63^27 = 4032\) \(k+s+g = 2(1 + 3^2 + 2^37)\) = 2b \(g/k = 2(2^27) = 2w\) ... similarly other 4 cases will also satisfy all 3 given constraints

ganeshie8 (ganeshie8):

wat do u mean by an uniq solution ha ?

ganeshie8 (ganeshie8):

unique age is not possible for anyone of kid/son/grandpa wid just those 3 constraints right ?

OpenStudy (dan815):

u want solution?

OpenStudy (dan815):

there is a unique solution

ganeshie8 (ganeshie8):

wit, clarify my question first

ganeshie8 (ganeshie8):

*Wait

OpenStudy (dan815):

ok lemme read

ganeshie8 (ganeshie8):

how are above two cases violating the given constraints ?

ganeshie8 (ganeshie8):

just read my recent reply where i marked the two lines in "red"

OpenStudy (dan815):

um lemme see there are certain constraints that come up when u see this equation

OpenStudy (dan815):

finding the constraints are actually part of the question

OpenStudy (dan815):

thats most of the hard work

OpenStudy (dan815):

like so u know that Z=12*sqrt(14/YW)

ganeshie8 (ganeshie8):

question boils down to only 3 constraints : 1) product of ages = 4032 2) sum of ages is even number 3) kids age evenly goes in grandpa's age

ganeshie8 (ganeshie8):

everything else can be derived based on above 3

OpenStudy (dan815):

lets say that k=P/(somefactor of 12)^2

OpenStudy (dan815):

no more constraints like

ganeshie8 (ganeshie8):

im still looking for explanation of how the two cases i gave dont fit the solution

OpenStudy (dan815):

id have to go into some of the deailts of the solution to show u some extra constraints

ganeshie8 (ganeshie8):

question boils down to only 3 constraints : 1) product of ages = 4032 2) sum of ages is even number 3) kids age evenly goes in grandpa's age

ganeshie8 (ganeshie8):

I thought, those are the only 3 constraints we need to satisfy ?

ganeshie8 (ganeshie8):

btw, the sqrt(2016/yw) thing is derived from above.. its not a new constraint :/

OpenStudy (dan815):

i know i was thinking about this too

OpenStudy (dan815):

i was having the same issue where u said there are other possible answers as long as yw is a factor of 12^2 and divides 14

ganeshie8 (ganeshie8):

so, the system is : ksg = 4032 k+s+g = 2b g/k = 2w k < s < g all naturals

ganeshie8 (ganeshie8):

can u derive any single other constraint, thats different from the above list /

ganeshie8 (ganeshie8):

?

ganeshie8 (ganeshie8):

nope, and the two cases i gave satisfy above constraints. so the solution is not unique QED.

ganeshie8 (ganeshie8):

^^thats my thinking...

OpenStudy (dan815):

hold on i argued this with my friend too

OpenStudy (dan815):

look at this

OpenStudy (dan815):

dont mind the pellets here and there he likes to swear -.- a lot

ganeshie8 (ganeshie8):

thats mouthful lol :/ il go thru this shortly, but u havent clarified my earlier question yet

OpenStudy (dan815):

so look for p such that 1< r <14 AND (14/r) is a non-prime integer

OpenStudy (dan815):

do u understand this statement here so look for p such that 1< r <14 AND (14/r) is a non-prime integer

OpenStudy (dan815):

honestly i dont know because like you i tried the other possibilies on the quiz

OpenStudy (dan815):

and it said it was wrong, they were looking for a unique solution and his solution was right

OpenStudy (dan815):

but then agian the question was taken down after a while, or i just couldnt find it again

OpenStudy (dan815):

im not sure if it was taken down because it wasnt a unique solution or i just cant find it -.-

ganeshie8 (ganeshie8):

there is a flaw in the solution

OpenStudy (dan815):

where

OpenStudy (dan815):

are u saying it because he said m is a ratio of grandpa to grandson

OpenStudy (dan815):

or "so look for p such that 1< r <14 AND (14/r) is a non-prime integerÈ

OpenStudy (dan815):

"so look for p such that 1< r <14 AND (14/r) is a non-prime integer"

ganeshie8 (ganeshie8):

nvm im still going thru it...

OpenStudy (dan815):

here was m solution i didnt fully understand his reasoning for that statement but i got same answer

OpenStudy (dan815):

but this has a flaw because u dont need a perfect square to come outta there it can be 1/2 1/3 or 1/6 aswell

OpenStudy (dan815):

not 1/4 since then y yould be odd and not 12 since then y<2

ganeshie8 (ganeshie8):

then we need (14/mz) = r (because y is an integer)

ganeshie8 (ganeshie8):

mz | 12^2*14 not just : mz | 14

OpenStudy (dan815):

yaa okkay makes sense so i guess the question was taken down because it was flawed

ganeshie8 (ganeshie8):

and when it divides, the result must be a perfect square

OpenStudy (dan815):

DANG it wasnt taken down before i lost some rating with a couple guesses

ganeshie8 (ganeshie8):

looking for mz | 14 gives unique solution when mz = 14, and y = 12 im not sure still i followed ur solution correctly.. uhmm

ganeshie8 (ganeshie8):

however there will be many other solutions when mz | 12^2 * 14 is a perfect square

ganeshie8 (ganeshie8):

nevertheless the problem is nice :) we can make the solution unique by changing the problem a bit i hope